OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, 10,5/3,5/18 , ...

OpenStudy (tkhunny):

Unfortunately, the squence is not well-defined. 60 * (1/6) = 10 10 * (1/6) = 10/6 = 5/3 (5/3) * (1/6) = 5/18 Okay, it APPEARS that it MAY be a simple geometric sequence and it clearly converges. However, we do NOT know that this finite pattern persists. The next term could be \(\sqrt{2}\) for all we know.